# On the Absolute-Value Integral of a Brownian Motion with Drift: Exact and Asymptotic Formulae

### v2

@@ 5 Asymptotic estimates
Since the numbers $-\alpha^{\prime}_{k-j}$, $k\geq 0$, are all positive and strictly increasing in $k$, this means that the G-function in (2) decays like $O(e^{2\sigma^{2}t^{3}(\alpha^{\prime}_{k-j})^{3}/(27x^{2})})$ as $1/x^{2}\rightarrow\infty$, and so the behavior of $f_{X_{t}}(x)$ for small $x$ is dominated by the leading-order term ($k=1$) in (2). In particular, letting $S$ be the $k$-series in (2), we have
@@
Under condition (i), let $\epsilon>0$ be such that the open ball with radius $\epsilon$ centered at $w$, $B_{\epsilon}(w)\subset D$. Consider the decomposition $\int_{D}=\int_{D\setminus B_{\epsilon}(w)}+\int_{B_{\epsilon}(w)}$. For the integral
$J_{0}(x)=\int_{D\setminus B_{\epsilon}(w)}g_{1}(w)e^{xg_{2}(w)-x^{2}g_{3}(w)}{\rm d}w,\quad x>0,$
@@
×$\qquad\times\Bigg{(}1+\sum^{\infty}_{m=1}\frac{1}{x^{m}m!}\bigg{(}\frac{1}{2}w^{\intercal}\mathrm{Hess}_{g_{2}}(\varpi_{2})w-\sum_{|a|=3}\bigg{(}\frac{\mathrm{D}^{a}g_{3}(w_{0})}{a!}+G_{a}\bigg{(}\frac{w}{x}+w_{0}\bigg{)}\bigg{)}w^{a}\bigg{)}^{m}\Bigg{)}{\rm d}w.$
@@
×$\qquad\times\bigg{(}\frac{1}{2}w^{\intercal}\mathrm{Hess}_{g_{2}}(\varpi_{2})w-\sum_{|a|=3}\bigg{(}\frac{\mathrm{D}^{a}g_{3}(w_{0})}{a!}+G_{a}\bigg{(}\frac{w}{x}+w_{0}\bigg{)}\bigg{)}w^{a}\bigg{)}^{m}{\rm d}w,$ (5.9)
@@
$f_{X_{t}}(x)=\sqrt{\frac{6}{\pi\sigma^{2}t^{3}}}\exp\bigg{(}-\frac{3\mu^{2}t}{8\sigma^{2}}-\frac{3x^{2}}{2\sigma^{2}t^{3}}\bigg{)}\cosh\frac{3\mu x}{2\sigma^{2}t}(1+O(x^{-1}))\quad\text{as }x\rightarrow\infty$ (5.10)
@@
$C_{d}$ =$=\{u:\mathrm{Re}u=d,\;|u|\leq R\},$
$C^{\delta}_{R}$ =$=\{u:\mathrm{Re}u\leq d,\;|u|=R,\;|\arg u-\theta|<\pi-\delta\},$
$C^{\delta}_{r}$ =$=\{u:\mathrm{Re}u\leq d,\;|u|=r,\;|\arg u-\theta|<\pi-\delta\},$
$C^{\pm}_{\delta}$ =$=\{u:\arg u=\theta\pm\pi\mp\delta,\;r\leq|u|\leq R\}.$
@@
$\int_{C^{+}_{0}}+\int_{C^{-}_{0}}=-\int^{\infty}_{0}\frac{e^{{\rm i}\theta}e^{-\rho e^{{\rm i}\theta}x}e^{-2{\rm i}\theta/3}}{(2\rho)^{2/3}(\rho e^{{\rm i}\theta}+{\rm i}\epsilon)}\big{(}e^{-2\pi{\rm i}/3}\psi_{+}(\rho,\theta)-e^{2\pi{\rm i}/3}\psi_{-}(\rho,\theta)\big{)}{\rm d}\rho,$ (5.15)
@@
$\qquad-{\rm i}(\bar{\mathrm{Ai}}(b(2\rho e^{{\rm i}(\theta-\pi)})^{-1/3})+\bar{\mathrm{Ai}}(b(2\rho e^{{\rm i}(\theta+\pi)})^{-1/3}))\mathrm{Bi}^{\prime}(\xi(\rho))\big{)},$
@@
which, for the same reason, exists as a double integral in the domain +×$\mathds{R}_{+}\times(c+{\rm i}\mathds{R})$.
@@
×$\qquad\times{\rm i}\bigg{(}A_{1}(\rho,\theta)+\int^{\xi(\rho,\theta)}_{0}A_{2}(\rho,\theta,z){\rm d}z\bigg{)}{\rm d}(\rho,\theta),$ (5.20)
@@
×$\quad\times\exp\bigg{(}\frac{{\rm i}\theta}{6}-\frac{2\sqrt{2}t^{3}e^{-{\rm i}\theta}(1+{\rm i}\tan\theta)^{3/2}}{9\varrho x}\bigg{(}\frac{9x^{2}}{2t^{4}}+\frac{b^{2}}{2}\bigg{)}^{3/2}\bigg{)}(1+O(x^{-2}))$
@@
×$\quad\times\big{(}e^{b\varPsi(\varrho,\theta)x/t}+e^{-b\varPsi(\varrho,\theta)x/t}\big{)}$
×$\quad\times\exp\bigg{(}\frac{{\rm i}b^{2}t\tan\theta}{2}-\frac{b^{2}te^{-{\rm i}\theta}(1+{\rm i}\tan\theta)^{3/2}}{2\varrho}-\frac{{\rm i}\theta}{2}-\frac{x^{2}}{t^{3}}\varPhi(\varrho,\theta)\bigg{)}{\rm d}(\varrho,\theta)\;(1+O(x^{-1})),$ (5.29)
@@
++×$\iint_{\mathds{R}_{++}\times(-\pi/2,\pi/2)}$ $=\frac{2\pi t^{3}e^{-b^{2}t/2}(e^{3bx/(2t)}+e^{-3bx/(2t)})e^{-3x^{2}/(2t^{3})}}{x^{2}\sqrt{\det\mathrm{Hess}_{\varPhi}(1,0)}}$
×$\quad\times\exp\bigg{(}\frac{b^{2}t(\nabla\varPsi)^{\intercal}(1,0)\mathrm{Hess}^{-1}_{\varPhi}(1,0)\nabla\varPsi(1,0)}{2}\bigg{)}(1+O(x^{-1}))$
@@
×$\qquad\times{\rm i}\bigg{(}A_{1}(\rho,\theta)+\int^{\xi(\rho,\theta)}_{0}A_{2}(\rho,\theta,z){\rm d}z\bigg{)}{\rm d}(\rho,\theta),\quad x>0,$
where $c>0$ is arbitrary, $b=\mu/\sigma$, $\xi(\rho,\theta)=2^{1/3}\rho^{-2/3}e^{-2{\rm i}\theta}(c+b^{2}/2)(1+{\rm i}\tan\theta)$, and $A_{1}$ and $A_{2}$ are as specified in (5).
@@ 6 Numerical experiments
$\alpha^{\prime}_{k}=-\bigg{(}\frac{3\pi(4k-1)}{8}\bigg{)}^{2/3}(1+O(k^{-2})),\quad\text{as }k\rightarrow\infty.$
@@
For the moments the formula (4.3) along with (4) and (4) is used for obtaining analytic expressions and their corresponding numerical values are all rounded to 10 significant digits. All implementations are done in ${}^{\circledR}$ ([38, Wolfram Research, Inc., 2023]) and results are reported in Table 1, [^11] Table 1 is the same as [30, Takács, 1993, Tab. 3]. Table 2, Table 3, and Table 4.
Importantly, for small to moderate values of $b=\mu/\sigma\neq 0$ it usually requires a high machine precision to correctly convert the exact values of some high-order moments to numerical values. Clearly, this challenge is caused by taking differences of large numbers subject to division by irrationals. Somewhat surprisingly, such numerical conversion can easily fail in many other programming languages, including ${}^{\circledR}$, at least in default settings – check, e.g., the last three rows in Table 2. To effectively avoid this issue, one may switch to the formula (4) with proper truncation of the infinite series, as discussed in Section 4, while the asymptotic estimate (5.31) can also come in handy when necessary.

### v4

@@ 5 Asymptotic estimates
Since the numbers $\alpha^{\prime}_{k-j}$, $k\geq 0$, are all positive and strictly increasing in $k$, this means that the G-function in (2) decays like $O(e^{2\sigma^{2}t^{3}(\alpha^{\prime}_{k-j})^{3}/(27x^{2})})$ as $1/x^{2}\rightarrow\infty$, and so the behavior of $f_{X_{t}}(x)$ for small $x$ is dominated by the leading-order term ($k=1$) in (2). In particular, letting $S$ be the $k$-series in (2), we have
@@
Under condition (i), let $\epsilon>0$ be such that the open ball with radius $\epsilon$ centered at $w_{0}$, $B_{\epsilon}(w_{0})\subset D$. Consider the decomposition $\int_{D}=\int_{D\setminus B_{\epsilon}(w_{0})}+\int_{B_{\epsilon}(w_{0})}$. For the integral
$J_{0}(x)=\int_{D\setminus B_{\epsilon}(w_{0})}g_{1}(w)e^{xg_{2}(w)-x^{2}g_{3}(w)}{\rm d}w,\quad x>0,$
@@
$\qquad\times\Bigg(1+\sum^{\infty}_{m=1}\frac{1}{x^{m}m!}\bigg(\frac{1}{2}w^{\intercal}\mathrm{Hess}_{g_{2}}(\varpi_{2})w-\sum_{|a|=3}\bigg(\frac{\mathrm{D}^{a}g_{3}(w_{0})}{a!}+G_{a}\bigg(\frac{w}{x}+w_{0}\bigg)\bigg)w^{a}\bigg)^{m}\Bigg){\rm d}w.$
@@
$\qquad\times\bigg(\frac{1}{2}w^{\intercal}\mathrm{Hess}_{g_{2}}(\varpi_{2})w-\sum_{|a|=3}\bigg(\frac{\mathrm{D}^{a}g_{3}(w_{0})}{a!}+G_{a}\bigg(\frac{w}{x}+w_{0}\bigg)\bigg)w^{a}\bigg)^{m}{\rm d}w,$ (5.9)
@@
$f_{X_{t}}(x)=\sqrt{\frac{6}{\pi\sigma^{2}t^{3}}}\exp\bigg(-\frac{3\mu^{2}t}{8\sigma^{2}}-\frac{3x^{2}}{2\sigma^{2}t^{3}}\bigg)\cosh\frac{3\mu x}{2\sigma^{2}t}(1+O(x^{-1})),\quad\text{as }x\rightarrow\infty$ (5.10)
@@
$C_{d}$ $=\{u:\mathrm{Re}u=d,\;|u|\leq R\},$
$C^{\delta}_{R}$ $=\{u:\mathrm{Re}u\leq d,\;|u|=R,\;|\arg u-\theta|<\pi-\delta\},$
$C^{\delta}_{r}$ $=\{u:\mathrm{Re}u\leq d,\;|u|=r,\;|\arg u-\theta|<\pi-\delta\},$
$C^{\pm}_{\delta}$ $=\{u:\arg u=\theta\pm\pi\mp\delta,\;r\leq|u|\leq R\}.$
@@
$\int_{C^{+}_{0}}+\int_{C^{-}_{0}}=\int^{\infty}_{0}\frac{e^{{\rm i}\theta}e^{-\rho e^{{\rm i}\theta}x}e^{-2{\rm i}\theta/3}}{(2\rho)^{2/3}(\rho e^{{\rm i}\theta}+{\rm i}\epsilon)}\big(e^{-2\pi{\rm i}/3}\psi_{+}(\rho,\theta)-e^{2\pi{\rm i}/3}\psi_{-}(\rho,\theta)\big){\rm d}\rho,$ (5.15)
@@
$\qquad-{\rm i}(\bar{\mathrm{Ai}}(b(2\rho e^{{\rm i}(\theta-\pi)})^{-1/3})+\bar{\mathrm{Ai}}(b(2\rho e^{{\rm i}(\theta+\pi)})^{-1/3}))\mathrm{Bi}^{\prime}(\xi(\rho))\big)\bigg),$
@@
which, for the same reason, exists as a double integral in the domain $\mathds{R}_{+}\times(c+{\rm i}\mathds{R})$.
@@
$\qquad\times{\rm i}\bigg(A_{1}(\rho,\theta)+\int^{\xi(\rho,\theta)}_{0}A_{2}(\rho,\theta,z){\rm d}z\bigg){\rm d}(\rho,\theta),$ (5.20)
@@
$\quad\times\exp\bigg(\frac{{\rm i}\theta}{6}-\frac{2\sqrt{2}t^{3}e^{-{\rm i}\theta}(1+{\rm i}\tan\theta)^{3/2}}{9\varrho x}\bigg(\frac{9x^{2}}{2t^{4}}+\frac{b^{2}}{2}\bigg)^{3/2}\bigg)(1+O(x^{-2}))$
@@
$\quad\times\big(e^{b\varPsi(\varrho,\theta)x/t}+e^{-b\varPsi(\varrho,\theta)x/t}\big)$
$\quad\times\exp\bigg(\frac{{\rm i}b^{2}t\tan\theta}{2}-\frac{b^{2}te^{-{\rm i}\theta}(1+{\rm i}\tan\theta)^{3/2}}{2\varrho}-\frac{{\rm i}\theta}{2}-\frac{x^{2}}{t^{3}}\varPhi(\varrho,\theta)\bigg){\rm d}(\varrho,\theta)\;(1+O(x^{-1})),$ (5.29)
@@
$\iint_{\mathds{R}_{++}\times(-\pi/2,\pi/2)}$ $=\frac{2\pi t^{3}e^{-b^{2}t/2}(e^{3bx/(2t)}+e^{-3bx/(2t)})e^{-3x^{2}/(2t^{3})}}{x^{2}\sqrt{\det\mathrm{Hess}_{\varPhi}(1,0)}}$
$\quad\times\exp\bigg(\frac{b^{2}t(\nabla\varPsi)^{\intercal}(1,0)\mathrm{Hess}^{-1}_{\varPhi}(1,0)\nabla\varPsi(1,0)}{2}\bigg)(1+O(x^{-1}))$
@@
$\qquad\times{\rm i}\bigg(A_{1}(\rho,\theta)+\int^{\xi(\rho,\theta)}_{0}A_{2}(\rho,\theta,z){\rm d}z\bigg){\rm d}(\rho,\theta),\quad x>0,$
where $c>0$ is arbitrary, $b=\mu/\sigma$, $\xi(\rho,\theta)=2^{1/3}\rho^{-2/3}e^{-2{\rm i}\theta/3}(c+b^{2}/2)(1+{\rm i}\tan\theta)$, and $A_{1}$ and $A_{2}$ are as specified in (5).
@@ 6 Numerical experiments
$\alpha^{\prime}_{k}=-\bigg(\frac{3\pi(4k-3)}{8}\bigg)^{2/3}(1+O(k^{-2})),\quad\text{as }k\rightarrow\infty.$
@@
For the moments the formula (4.3) along with (4) and (4) is used for obtaining analytic expressions and their corresponding numerical values are all rounded to 10 significant digits. All implementations are done in Mathematica® ([38, Wolfram Research, Inc., 2023]) and results are reported in Table 1, [^11] Table 1 is the same as [30, Takács, 1993, Tab. 3]. Table 2, Table 3, and Table 4.
Importantly, for small to moderate values of $b=\mu/\sigma\neq 0$ it usually requires a high machine precision to correctly convert the exact values of some high-order moments to numerical values. Clearly, this challenge is caused by taking differences of large numbers subject to division by irrationals. Somewhat surprisingly, such numerical conversion can easily fail in many other programming languages, including Matlab®, at least in default settings – check, e.g., the last three rows in Table 2. To effectively avoid this issue, one may switch to the formula (4) with proper truncation of the infinite series, as discussed in Section 4, while the asymptotic estimate (5.31) can also come in handy when necessary.